If a square has an area of 200 yd. squared. What will the area of the square become if you double the length of the sides.
What i have tried:
Figuring out the sides now when it is 200, my problem is that i can't figure out those sides.
Thank you!

Answer by checkley77(12844)

X^2=200
X=SQRT200
X=14.1421335 YDS. IS THE LENGTH OF THE SIDES OF THIS SQUARE.
2*14.142135=28.28427
28.28427^2=800 YDS.^2 IS THE NEW AREA OR THE AREA IS MULTIPLIED BY 4.